# Elementary Analysis of Policy Gradient Methods

### v1

@@ 5 Softmax Natural Policy Gradient
$\log Z_{s}^{k}=\mathrm{KL}(\pi^{k}_{s}\|\pi^{k+1}_{s}).$
Taking expectation with respect to $\pi^{k+1}_{s}$ yields
@@
Taking expectation with respect to $\pi^{*}_{s}$ yields
@@
$=(1-b_{s}^{k})\,\mathbb{E}_{s\sim\xi^{k}(\cdot|s)}\left[\exp\left(\eta\,A^{k}_{s,a}\right)\right]+b_{s}^{k}\,\mathbb{E}_{a^{\prime}\sim\xi^{k}_{-}(\cdot|s)}\left[\exp\left(\eta\,A^{k}_{s,a^{\prime}}\right)\right]$
@@
Together with Theorem 2.13, the improvement lower bounds I and II imply the global linear convergence of softmax NPG. ∎
Note that the following dynamic convergence result solely based on the improvement bound I holds for all the $k$,
@@
Next we consider a bandit example where the rate which shows that this dynamic is indeed tight, i.e., the equality can be achieved.
@@
Moreover,
$T(\sigma)\leq\frac{2}{\sigma\,\tilde{\rho}}\left[\frac{1}{\eta\,(1-\gamma)}\cdot\mathrm{KL}_{d_{\rho}^{*}}(\pi^{*}\|\pi^{0})+\frac{1}{(1-\gamma)^{2}}\right].$
@@
Let $s_{0}=\arg\min_{s}\{\max_{a}Q^{*}(s,a)-\max_{a\not\in\mathcal{A}_{s}^{*}}Q^{*}(s,a)\}$ and $a_{0}\in\arg\max_{s_{0}}Q^{*}(s_{0},a)$. According to the definition of $\Delta$, one has
@@
and the first claim follows immediately. Since
$\|A^{*}-A^{k}\|_{\infty}\leq\|V^{*}-V^{k}\|_{\infty}\leq\frac{1}{\tilde{\rho}}\left(V^{*}(\rho)-V^{k}(\rho)\right),$
the second claim follows directly from the sublinear convergence result of softmax NPG. ∎
@@ 6.1 Preliminaries
where $r(s,a)\in[0,1]$ and $\mathcal{H}(p)=-\sum_{a}p_{a}\log p_{a}$ defines the entropy of a distribution. Moreover, the action value and advantage functions are defined as
@@
$V_{\tau}^{\pi}(s)\in\left[0,\frac{1+\tau\log|\mathcal{A}|}{1-\gamma}\right]\quad\mbox{and}\quad Q_{\tau}^{\pi}(s,a)\in\left[0,\frac{1+\gamma\,\tau\log|\mathcal{A}|}{1-\gamma}\right].$
@@ 6.2 Entropy Softmax PG
where the second line follows from $\eta\,\tilde{\mu}\leq\eta_{s}\leq\eta/(1-\gamma)$. The proof of the first claim is now complete, and the the second claim can be verified directly. ∎
@@
where the last inequality utilizes $\tau/c<2(1-\gamma)$ for $|{\cal A}|>1$. By the monotonically decreasing property of the function $\exp\left(-\frac{2\,x}{1-\gamma}c\right)-\frac{\tau\,x}{2(1-\gamma)}$, it can be concluded that $2\alpha<2\alpha^{\prime}<\beta$, i.e. we provide a wider range of converging step sizes for entropy softmax PG to converge linearly.
@@
$\leq\frac{1+\tau\log|{\cal A}|}{1-\gamma}\cdot\exp\left(-k\cdot(1-\gamma)\,\alpha\,\tilde{\mu}\,\tau\,\kappa^{2}\right)$
$\leq\frac{1+\tau\log|{\cal A}|}{1-\gamma}\cdot\exp\left(-(k-1)\cdot(1-\gamma)\,\alpha\,\tilde{\mu}\,\tau,\kappa^{2}\right),$
@@
For softmax PG, we have shown the convergence of algorithm for any constant step size (though with sublinear convergence rate). For entropy softmax PG, the convergence can only be guaranteed for the step size over a finite interval. Thus, one may wonder whether the convergence of entropy can be established for any constant step size. To investigate this issue, numerical test has been conducted on a random MDP, and the result suggests entropy softmax PG does not converge for a large step size, see Figure 1. In the test, the rewards for the state-action pairs and the entries of the transition probability matrix are generated through a uniform distribution over $\left(0,1\right)$. We then re-scale each row of the transition probability matrix to obtain a stochastic matrix.
@@ 6.3 Entropy Softmax NPG
The linear convergence of entropy regularized softmax NPG in terms of infinity norm has been established in [5, Theorem 1].
@@
the linear convergence of entropy regularized state value function follows immediately:
@@
Noting that $\rho$ can be an arbitrary initial distribution over $\mathcal{S}$, the proof is complete. ∎
@@
This basically means soft PI will convergence at least linearly with rate $\gamma$ to achieve the accuracy $\frac{2\tau(1-\gamma)}{\gamma^{2}}$, and then converges quadratically. Therefore, the result in Theorem 6.13 encodes both the information of the linear convergence phase and the quadratic convergence phase simultaneously.
@@
Since entropy softmax NPG tends to become soft PI when $\eta\rightarrow\infty$, it is natural to anticipate the convergence of entropy softmax NPG becomes faster as $\eta$ increases, at least locally due to the above local quadratic convergence of soft PI. However, the global rate in (6.11) converges to the fixed $\gamma^{2}$ as $\eta\rightarrow\infty$, which cannot reflects this intuition. At the end of this section, we are going to establish a local and tight convergence rate for entropy softmax NPG which can be arbitrarily small as $\eta$ increases. To this end, we first list two useful lemmas.
@@
$\lim_{k\rightarrow\infty}\frac{\mathrm{KL}(\pi^{k}\|\pi^{k+1})}{\mathrm{KL}(\pi^{{k+1}}\|\pi^{{k}})}=1\quad\mbox{and}\quad\lim_{k\rightarrow\infty}\frac{\mathrm{KL}(\pi^{k}\|\pi^{*})}{\mathrm{KL}(\pi^{*}\|\pi^{k})}=1.$
@@
These properties will be used in the establishment of the fast convergence rate of entropy softmax NPG. The improvement lower and upper bounds of $\mathcal{L}_{k}^{k+1}$ in terms of $\mathcal{L}_{k}^{*}$ (see (6.2) and (6.3) for their definitions) are firstly established.
@@
Theorem 6.17 implies that the local convergence rate of entropy softmax NPG is about ${1}/{(\eta\tau+1)^{2}}$, which becomes faster as $\eta\rightarrow\infty$. This on the other side suggests the local quadratic convergence of soft PI. Numerical results on random MDP demonstrates that the theoretical rate desirably matches the empirical one, see Figure 2.
@@ 7 Conclusion
In this paper, we study the convergence of a family of basic policy gradient methods for RL (including PPG, softmax PG, softmax NPG, entropy softmax PG, entropy softmax NPG, and soft PI) in a comprehensive way. New convergence results have been obtained for them based on elementary analysis techniques. There are several lines of direction for future work. For instance, it is interesting to see whether the analysis techniques can be used to handle the stochastic case where the policy gradient is evaluated through samples or the case in which policy parameterization involves function approximation. It is also likely to extend the analysis study to other scenarios in addition to the discounted MDP setting, such as the average reward MPD setting or the risk sensitive RL problem.
@@
As $f$ and $g$ are monotonically increasing we have
$\forall x,y\in\mathbb{R},\quad(f(x)-f(y))(g(x)-g(y))\geq 0,$
hence $\mathbb{E}[(f(X)-f(Y))(g(X)-g(Y))]\geq 0$ for arbitrary random variables $X,Y$. Consider $Y$ as an i.i.d. copy of $X$, we get $\mathrm{Cov}(f(X),g(X))\geq 0$ from Lemma 2.6. ∎
@@
$\geq\tilde{\rho}\,\mathbb{E}_{s\sim d_{\rho}^{*}}\left[\max_{a}A^{k}(s,a)\right],$
@@
Then $V^{k+1}(s)-V^{k}(s)\geq 0$ follows by the performance difference lemma and $Q^{k+1}(s,a)-Q^{k}(s,a)$ follows since $Q^{t}(s,a)=r(s,a)+\gamma\mathbb{E}_{s^{\prime}\sim P(\cdot|s,a)}\left[V^{t}(s^{\prime})\right]$.
@@
$\geq\frac{1}{1-\gamma}\sum_{s}d^{k+1}_{\mu}(s)\frac{1}{|\mathcal{A}|}\left(\max_{a}|\hat{A}^{k}(s,a)|\right)\left(1-\exp\left(-\alpha\,\tilde{\mu}\max_{a}|\hat{A}^{k}(s,a)|\right)\right)$
@@
Thus the second condition is met due to the expression of the gradient. ∎
@@
    For all $s,a$, the limit of the policy sequence $\{\pi^{k}(a|s)\}$ exists, i.e. there exists only one limit point.
@@
$\forall\,s\in{\cal S}:\quad{\cal T}_{\tau}^{k+1}V^{k}_{\tau}(s)-V^{k}_{\tau}(s)\geq 0,$
Hence, by the performance difference lemma (Lemma 6.1), $\left\{V^{k}_{\tau}(s)\right\}$ is a monotonically increasing sequence. Noting that $|V^{k}_{\tau}(s)|\leq(1+\tau\log|{\cal A}|)/(1-\gamma)$ is bounded, the limit exists, denoted
@@
By the relation of $Q^{k}_{\tau}(s,a)=r(s,a)+\gamma\mathbb{E}_{s^{\prime}\sim P(\cdot|s,a)}\left[V^{k}_{\tau}(s^{\prime})\right]$, we know that $Q^{\infty}_{\tau}(s,a)$ also exists, and
@@
where $C(\eta):=\eta\,\tilde{\mu}\left[\exp\left(-\frac{2\eta(1+\tau\log|{\cal A}|)}{(1-\gamma)^{2}}\right)-\frac{\tau\eta}{2(1-\gamma)}\right]$ is the constant given in Lemma 6.7. Noting that $d^{k+1}_{\mu}(s)\geq(1-\gamma)\tilde{\mu}>0$ and $C(\eta)>0$, together with $\lim_{k}\left(V^{k+1}(\mu)-V^{k}(\mu)\right)=0$ we get
@@
Note that $p(a)$ satisfies $\left[Q^{\infty}_{\tau}(s,a)-V^{\infty}_{\tau}(s)-\tau\log p(a)\right]=0$. For the case that $p(a)>1$, we have $Q^{\infty}_{\tau}(s,a)>V^{\infty}_{\tau}(s)$. Therefore there exists a constant $c_{0}>0$ and a finite time $T_{0}$ such that $Q^{k}_{\tau}(s,a_{+})-V^{k}_{\tau}(s)\geq c_{0}$ for all $k>T_{0}$, hence $Q^{k}_{\tau}(s,a_{+})-V^{k}_{\tau}(s)-\tau\log\pi^{k}(a_{+}|s)\geq c_{0}$ for all $k>T_{0}$. It follows from (D.2) that $\lim_{k\to\infty}\pi^{k}(a_{+}|s)=0$, i.e. only one limit point for $\{\pi^{k}(a|s)\}$.
For the case that $p(a)\in(0,1]$, we next show that (D.2) implies that $\pi^{k}(a|s)$ has two limit points $\{0,\,p(a)\}$, denoted $\pi^{k}(a|s)\to\{0,\,p(a)\}$. First we have
@@
$c_{k}\rightarrow 0,\quad\|\pi_{\xi_{k}}-\pi_{\theta^{*}}\|_{2}\rightarrow 0,\quad\|\pi_{\hat{\xi}_{k}}-\pi_{\theta^{*}}\|_{2}\rightarrow 0,\quad\mathrm{and}\|\pi_{\hat{\xi}_{k}}-\pi_{\theta^{*}}\|_{\infty}\rightarrow 0.$

### v2

@@ 5 Softmax Natural Policy Gradient
$\log Z_{s}^{k}=\mathrm{KL}(\pi^{k}_{s}\|\pi^{k+1}_{s}),$
while taking expectation with respect to $\pi^{k+1}_{s}$ yields
@@
Moreover, taking expectation with respect to $\pi^{*}_{s}$ on both sides of (5.2) yields
@@
$=(1-b_{s}^{k})\,\mathbb{E}_{a\sim\xi^{k}(\cdot|s)}\left[\exp\left(\eta\,A^{k}_{s,a}\right)\right]+b_{s}^{k}\,\mathbb{E}_{a^{\prime}\sim\xi^{k}_{-}(\cdot|s)}\left[\exp\left(\eta\,A^{k}_{s,a^{\prime}}\right)\right]$
@@
Together with Theorem 2.13, the improvement lower bounds I and II imply the global linear convergence of softmax NPG. Note that $\inf_{t}C^{t}$ exists since $b_{s}^{t}\rightarrow 0$ as softmax NPG converges. ∎
It is not hard to see that the following dynamic convergence result solely based on the improvement bound I holds for all the $k$,
@@
Next we consider a bandit example which shows that the above dynamic is indeed tight, i.e., the equality can be achieved.
@@
Let $s_{0}=\arg\min_{s}\{\max_{a}Q^{*}(s,a)-\max_{a\not\in\mathcal{A}_{s}^{*}}Q^{*}(s,a)\}$ and $a_{0}\in\arg\max_{a}Q^{*}(s_{0},a)$. According to the definition of $\Delta$, one has
@@
and the claim follows immediately. ∎
@@ 6.1 Preliminaries
where $r(s,a)\in[0,1]$ and $\mathcal{H}(p)=-\sum_{a}p_{a}\log p_{a}$ defines the entropy of a distribution. Moreover, the action value and the advantage function are defined as
@@
[…]
@@ 6.2 Entropy Softmax PG
where the second line follows from $\eta\,\tilde{\mu}\leq\eta_{s}\leq\eta/(1-\gamma)$. The proof of the first claim is now complete, and the second claim can be verified directly. ∎
@@
where the last inequality utilizes $\tau/c<2(1-\gamma)$ for $|{\cal A}|>1$. By the monotonically decreasing property of the function $\exp\left(-\frac{2\,x}{1-\gamma}c\right)-\frac{\tau\,x}{2(1-\gamma)}$, it can be concluded that $2\alpha<2\alpha^{\prime}<\beta$, i.e. we provide a wider range of step sizes for entropy softmax PG to converge linearly.
@@
$\leq\frac{1+\tau\log|{\cal A}|}{1-\gamma}\cdot\exp\left(-k\cdot(1-\gamma)\,\alpha\,\tilde{\mu}\,\tau\,\kappa^{2}\right),$
@@
For softmax PG, we have established its convergence for any constant step size (though with sublinear convergence rate). For entropy softmax PG, the convergence can only be guaranteed for the step size over a finite interval. Thus, one may wonder whether the convergence of entropy softmax PG can be established for any constant step size. To investigate this issue, numerical test has been conducted on a random MDP, and the result suggests entropy softmax PG does not converge for a large step size, see Figure 1. In the test, the rewards for the state-action pairs and the entries of the transition probability matrix are generated through a uniform distribution over $\left(0,1\right)$. We then re-scale each row of the transition probability matrix to obtain a stochastic matrix.
@@ 6.3 Entropy Softmax NPG
The linear convergence of entropy softmax NPG in terms of infinity norm has been established in [5, Theorem 1].
@@
the linear convergence in terms of the entropy regularized state value follows immediately:
@@
Noting that $\rho$ can be an arbitrary distribution over $\mathcal{S}$, the proof is complete. ∎
@@
This basically means soft PI will convergence at least linearly with rate $\gamma$ to achieve the accuracy $\frac{2\tau(1-\gamma)}{\gamma^{2}}$, and then converges quadratically. Therefore, the result in Theorem 6.13 encodes the information of the linear convergence phase and the quadratic convergence phase simultaneously.
@@
Since entropy softmax NPG tends to become soft PI when $\eta\rightarrow\infty$, it is natural to anticipate the convergence of entropy softmax NPG becomes faster (at least locally) as $\eta$ increases due to the above local quadratic convergence of soft PI. However, the global rate in (6.11) converges to the fixed $\gamma^{2}$ as $\eta\rightarrow\infty$, which cannot reflects this intuition. At the end of this section, we are going to establish a local and tight convergence rate for entropy softmax NPG which can be arbitrarily small as $\eta$ increases. To this end, we first list two useful lemmas.
@@
[…]
@@
These properties will be used in the establishment of the fast local convergence rate of entropy softmax NPG. The improvement lower and upper bounds of $\mathcal{L}_{k}^{k+1}$ in terms of $\mathcal{L}_{k}^{*}$ (see (6.2) and (6.3) for their definitions) are firstly established.
@@
Theorem 6.17 implies that the local convergence rate of entropy softmax NPG is about ${1}/{(\eta\tau+1)^{2}}$, which becomes faster as $\eta\rightarrow\infty$. This on the other side suggests the local quadratic convergence of soft PI. Numerical results on a random MDP demonstrates that the theoretical rate desirably matches the empirical one, see Figure 2.
@@ 7 Conclusion
In this paper, we comprehensively study the convergence of a family of basic policy gradient methods for RL, including PPG, softmax PG, softmax NPG, entropy softmax PG, entropy softmax NPG, and soft PI. New convergence results have been obtained for them based on elementary analysis techniques. There are several lines of direction for future work. For instance, it is interesting to see whether the analysis techniques can be used to handle the stochastic case where the policy gradient is evaluated through samples or the case in which policy parameterization involves function approximation. It is also likely to extend the analysis study to other scenarios in addition to the discounted MDP setting, such as the average reward MPD setting or the risk sensitive RL problem.
@@
As $f$ and $g$ are both monotonically increasing, we have
$\forall\,x,y\in\mathbb{R},\quad(f(x)-f(y))(g(x)-g(y))\geq 0.$
It follows that $\mathbb{E}[(f(X)-f(Y))(g(X)-g(Y))]\geq 0$ for arbitrary random variables $X,Y$. Considering $Y$ as an i.i.d. copy of $X$, we get $\mathrm{Cov}(f(X),g(X))\geq 0$ from Lemma 2.5. ∎
@@
$\geq\tilde{\rho}\,\mathbb{E}_{s\sim d_{\rho}^{*}}\left[\max_{a}A^{k}(s,a)\right].$
@@
Then $V^{k+1}(s)-V^{k}(s)\geq 0$ follows by the performance difference lemma and $Q^{k+1}(s,a)-Q^{k}(s,a)\geq 0$ follows from $Q^{t}(s,a)=r(s,a)+\gamma\mathbb{E}_{s^{\prime}\sim P(\cdot|s,a)}\left[V^{t}(s^{\prime})\right]$.
@@
$\geq\frac{1}{1-\gamma}\sum_{s}d^{k+1}_{\mu}(s)\frac{1}{|\mathcal{A}|}\left(\max_{a}|\hat{A}^{k}(s,a)|\right)\left(1-\exp\left(-\alpha\,\tilde{\mu}\max_{a}|\hat{A}^{k}(s,a)|\right)\right).$
@@
Therefore the second condition further holds due to the expression of the gradient. ∎
@@
    For all $s,a$, the limit of the policy sequence $\{\pi^{k}(a|s)\}$ indeed exists, namely there exists only one limit point.
@@
$\forall\,s\in{\cal S}:\quad{\cal T}_{\tau}^{k+1}V^{k}_{\tau}(s)-V^{k}_{\tau}(s)\geq 0.$
Hence, by the performance difference lemma (Lemma 6.1), $\left\{V^{k}_{\tau}(s)\right\}$ is a monotonically increasing sequence. Noting that $|V^{k}_{\tau}(s)|\leq(1+\tau\log|{\cal A}|)/(1-\gamma)$ is bounded, the limit exists,
@@
By the relation of $Q^{k}_{\tau}(s,a)=r(s,a)+\gamma\mathbb{E}_{s^{\prime}\sim P(\cdot|s,a)}\left[V^{k}_{\tau}(s^{\prime})\right]$, we know that $Q^{\infty}_{\tau}(s,a):=\lim_{k\rightarrow\infty}Q^{k}_{\tau}(s,a)$ also exists, and
@@
where $C(\eta):=\eta\,\tilde{\mu}\left[\exp\left(-\frac{2\eta(1+\tau\log|{\cal A}|)}{(1-\gamma)^{2}}\right)-\frac{\tau\eta}{2(1-\gamma)}\right]$ is the constant given in Lemma 6.7. Noting that $d^{k+1}_{\mu}(s)\geq(1-\gamma)\tilde{\mu}>0$ and $C(\eta)>0$, together with $\lim_{k}\left(V^{k+1}(\mu)-V^{k}(\mu)\right)=0$, we get
@@
Note that $p(a)$ satisfies $Q^{\infty}_{\tau}(s,a)-V^{\infty}_{\tau}(s)-\tau\log p(a)=0$.
For the case that $p(a)>1$, we have $Q^{\infty}_{\tau}(s,a)>V^{\infty}_{\tau}(s)$. Therefore there exists a constant $c_{0}>0$ and a finite time $T_{0}$ such that $Q^{k}_{\tau}(s,a)-V^{k}_{\tau}(s)\geq c_{0}$ for all $k>T_{0}$, so $Q^{k}_{\tau}(s,a)-V^{k}_{\tau}(s)-\tau\log\pi^{k}(a|s)\geq c_{0}$ for all $k>T_{0}$. It follows from (D.2) that $\lim_{k\to\infty}\pi^{k}(a|s)=0$, i.e. only one limit point for $\{\pi^{k}(a|s)\}$.
For the case that $p(a)\in(0,1]$, we next show that (D.2) implies that $\pi^{k}(a|s)$ has at most two limit points $\{0,\,p(a)\}$, denoted $\pi^{k}(a|s)\to\{0,\,p(a)\}$. First we have
@@
$c_{k}\rightarrow 0,\quad\|\pi_{\xi_{k}}-\pi_{\theta^{*}}\|_{2}\rightarrow 0,\quad\|\pi_{\hat{\xi}_{k}}-\pi_{\theta^{*}}\|_{2}\rightarrow 0,\quad\mbox{and }\|\pi_{\hat{\xi}_{k}}-\pi_{\theta^{*}}\|_{\infty}\rightarrow 0.$